Given that Eo for the reaction below is +1.51 V,


(a)       calculate the reduction potential, E, in a solution of pH 2.5 and in which the ratio [Mn2+]:[MnO4-] = 1: 100.

(b)       Show that E = +1:25 V when pH ¼ 3.0 and the ratio [Mn2+]:[MnO4-] =1:100.

(c)       For a ratio [Mn2+]:[MnO4-]=1000 : 1, what must the pH of the solution be to give a value of E = +1.45 V?

(d)       For a ratio [[Mn2+]: [MnO4-]=1 : 100, determine E in a solution of pH 1.8.



1
Expert's answer
2021-07-13T02:42:56-0400

we have,

MnO4​ + 8H+ + 5e→ Mn2+ + 4H2​O


we know that


"E=E^o-{0.059 \over 5}log{[Mn^{2+}]\over [MnO_4^{-}][H^{+}]^8}"


putting Eo=+1.51


"E=1.51-{0.059 \over 5}log{[Mn^{2+}]\over [MnO_4^{-}][H^{+}]^8}" .............................1


a)

pH= 2.5 taking antilog we get

[H+]=3.16 x 10-3


[Mn2+]:[MnO4-] = 1: 100.


putting in !1 we get,


"E=1.51-{0.059 \over 5}log{[1]\over [100][3.16 \times10^{-3}]^8}"


"E=1.51-{0.06\over5} \times18"


"E=1.294"


b) To show E=1.25 V

pH= 3 taking antilog we get

[H+]= "10^{-3}"


[Mn2+]:[MnO4-] = 1: 100.


putting in !1 we get,


"E=1.51-{0.059 \over 5}log{[1]\over [100][10^{-3}]^8}"


"E=1.51-{0.059 \over 5}\times22"


"E=1.25 V" APPROX


c)

pH= ?

[H+]=?

E = +1.45 V

[Mn2+]:[MnO4-] = 1000 : 1


putting in !1 we get,


"1.45=1.51-{0.059 \over 5}log{[1000]\over [1][H^{+}]^8}"


"0.06={0.06\over5}log{[1000]\over [1][H^{+}]^8}"

"10^5={1\over 1000 H^+}"


"[H^+]=10^{-7}"


d)

pH= 1.8 taking antilog we get

[H+]=1.5 X 10-2


[Mn2+]:[MnO4-] = 1: 100.


putting in !1 we get,


"E=1.51-{0.059 \over 5}log{[1]\over [100][1.5\times 10^{-2}]^8}"


"E=1.51-{0.059\over 5}{(14-1.6)}"


"E=1.361" V